OpenStudy (anonymous):

What is the missing number in the arithmetic sequence -11, 5, __, 37, ...?

OpenStudy (anonymous):

Table[ 16n-27, { n, 6 } ] -> { -11, 5, 21, 37, 53, 69 }

OpenStudy (lynfran):

-11+16=5 5+16=21 21+16=37 37+16=53 53+16=69